OpenStudy (anonymous):

A number is multiplied by 10. Then 32 is added to the product. This sum is then divided by 4. The answer is 28. What was the original number?

OpenStudy (anonymous):

@JustDoubleChecking

OpenStudy (texaschic101):

(10x + 32) / 4 = 28 multiply both sides by 4

OpenStudy (texaschic101):

notice that when u do this, the 4's on the left side cancel

OpenStudy (texaschic101):

leaving u with... 10x + 32 = 28 * 4 can u solve for x now ?

OpenStudy (anonymous):

I know 7 x 4 = 28

OpenStudy (anonymous):

suppose to work backwards to find out the beginning #

OpenStudy (texaschic101):

10x + 32 = 112 10x = 112 - 32 10x = 80 x = 80/10 x = 8

OpenStudy (texaschic101):

u have any q's about what i did ?

OpenStudy (anonymous):

original number is 8

OpenStudy (texaschic101):

yes

OpenStudy (anonymous):

thanks, I understand

OpenStudy (texaschic101):

yw